### v2

# SEED: Customize Large Language Models with Sample-Efficient Adaptation for Code Generation

Abstract
Although Large Language Models (LLMs) have made significant progress in code generation, they still struggle with code generation tasks in specific scenarios. These scenarios usually necessitate the adaptation of LLMs to fulfill specific needs, but the limited training samples available in practice lead to poor code generation performance. Therefore, how to effectively adapt LLMs to new scenarios with few training samples is a major challenge for current code generation. In this paper, we propose a novel adaptation approach named SEED, which stands for Sample-Efficient adaptation with Error-Driven learning for code generation. SEED leverages the errors made by LLMs as learning opportunities, using error revision to overcome its own shortcomings, thus achieving efficient learning. Specifically, SEED involves identifying error code generated by LLMs, employing Self-Revise for code revision, optimizing the model with revised code, and iteratively adapting the process for continuous improvement. Experimental results show that, compared to other mainstream fine-tuning approaches, SEED achieves superior performance with few training samples, showing an average relative improvement of 54.7% in Pass@1 on multiple code generation benchmarks. We also validate the effectiveness of Self-Revise, which generates revised code that optimizes the model more efficiently compared to the code samples from datasets. Moreover, SEED consistently demonstrates strong performance across various LLMs, underscoring its generalizability.
Index Terms:
Code Generation, Sample-Efficient Adaptation, Large Language Model.
I Introduction
Code generation is an important technology that can improve the efficiency and quality of software development. Given the human requirement expressed in natural language, code generation allows machines to generate executable programs that satisfy this requirement. Code generation has been a research hot topic in the fields of artificial intelligence, software engineering, and natural language processing. Recently, code generation technologies have made significant advancements in both academia and industry [1, 2, 3]. In particular, large language models (LLMs) demonstrate great potential in code generation tasks [4, 5, 6, 7, 8, 9]. However, LLMs still face significant challenges in code generation for some specific scenarios or domains [10, 11].
For specific code generation scenarios, fine-tuning is an essential adaptation approach to ensure LLMs fulfill particular needs [12, 13, 14, 15]. However, in these specific scenarios, it is difficult to obtain sufficient training samples for fine-tuning LLMs, due to common reasons such as industry secrecy and scarcity of resources. For example, in safety-critical scenarios like aerospace, medical devices, transportation, and financial industries, the generated code must adhere to specific security specifications, and accessing relevant data is often extremely difficult due to high confidentiality and strict access control. Under the circumstance of limited data, mainstream fine-tuning approaches might not enable LLMs to achieve the desired code generation performance and may even lead to a degradation in model performance [16, 17, 18], as shown in Figure 1. Consequently, how to effectively adapt LLMs to specific scenarios with limited data available is a major challenge for code generation in practice.
Figure 1: The performance of direct generation, fine-tuning, and our proposed SEED on MBPP dataset with few training samples. The numbers on the bars indicate the training samples used by different approaches.
The mainstream fine-tuning approaches use a large number of samples gathered under specific scenarios for training [19]. They enable model to exhaustively learn the features present in these samples and thus adapt to the specific scenarios. However, they have two disadvantages. First, compelling LLMs to relearn the entire code data of these new scenarios is inefficient. Considering that LLMs are pre-trained on large-scale and diverse data, it’s reasonably assumed that they possess a certain level of general knowledge, lacking only particular information for application in specific scenarios. Second, when faced with insufficient data volume or data drift, model may learn certain undesirable features (such as inaccurate or irrelevant programming knowledge and patterns), thereby affecting its learning efficiency and negatively impacting its final performance.
To overcome the disadvantages of these mainstream fine-tuning approaches, we take inspiration from the error-driven learning observed in humans. First, error-driven learning requires learners to identify their errors through testing. It helps learners to identify what they have mastered and what they still need to learn, allowing them to narrow the scope of learning and avoid wasting efforts on irrelevancies. Second, through error revision, learners can understand their deficiencies and make targeted improvements, thus enhancing learning efficiency and effectiveness. This motivates us to explore approaches to achieve sample-efficient adaptation of LLMs for code generation guided by error-driven learning.
In this paper, we propose SEED, a Sample-Efficient adaptation based on Error-Driven learning for code generation. SEED aims to alleviate the problem of poor code generation performance of fine-tuning LLMs in scenarios with few training samples. Following the error-driven learning, our approach proceeds in four steps: ❶ Error Code Collection. We identify and collect error codes generated by LLMs, aiming to mine the weaknesses of LLMs. ❷ Automatic Code Revision. To obtain revisions of error codes in a low-cost way, we design Self-Revise to realize automatic revision leveraging information in the original dataset and code execution feedback. ❸ Model Optimization. We optimize the LLMs using the revised code, making them focus on learning the revision of these critical errors, thereby improving the learning efficiency of LLMs. ❹ Iterative Adaptation. We adopt an iterative strategy, which involves repeating the preceding three steps, to continuously optimize and improve the performance of LLMs. Extensive experimental results demonstrate the superiority and generalizability of SEED in the sample-efficient adaptation of LLMs for specific code generation scenarios.
To summarize, the main contributions of this paper are:
  • We introduce error-driven learning for LLMs. It utilizes revisions of model’s erroneous outputs for training and has higher training efficiency compared to using samples in the dataset.
  • We propose a sample-efficient adaptation approach of LLMs for code generation, named SEED. Based on the principle of error-driven learning, SEED continuously optimizes model by identifying and revising error codes generated by LLMs, enabling effective adaptation to specific code generation scenarios with limited samples.
  • SEED outperforms other mainstream fine-tuning approaches on three public code generation benchmarks across various LLMs with few training samples.
II Motivation Example
Figure 2: A motivation example of SEED.
Figure 3: An overview of the proposed SEED and its differences from traditional fine-tuning approaches.
Aligning LLMs with specific scenarios and addressing their unique challenges by learning samples in the dataset is difficult, especially when training samples are limited. We present a motivation example in Figure 2 to clarify the advantages of using error-driven learning in the LLMs adaptation process of code generation.
[…]
Further, we explore the effectiveness of adopting error-driven learning from the perspective of model optimization. We consider the potential significant discrepancy between the model-generated output and the sample in the dataset. By learning the revisions of the model’s erroneous outputs, we can find more effective navigation in the optimization process. This might provide a shorter, smoother path to a good local minimum compared to learning from samples in the dataset, rather than attempting to direct it toward a distant area that may not align well with its existing knowledge or biases. We conduct the statistical analysis of the discrepancies in the model’s latent representations [^2] Specifically, on MBPP dataset, we obtain erroneous outputs of CodeGen-2B, revisions of the outputs, and samples in MBPP. We concatenate the requirements with their code, input them into CodeGen-2B, and extract the hidden representations from the model’s final layer. Then, we compute the Euclidean distances within the model’s representational space to quantify the disparities between these three elements.. The findings reveal that the average distance between the model’s erroneous outputs and the dataset’s samples is 12.35, whereas the average distance between the erroneous outputs and their revisions is significantly lower, at 6.39. These experimental results suggest that within the model’s representation space, revised codes are closer and similar to the erroneous output codes than the original code samples. This evidence lends support to our hypothesis of why the error-driven learning method is more efficient.
Therefore, our work is determined to explore the use of error-driven learning to achieve a sample-efficient adaptation approach, aimed at enhancing the performance of LLMs in specific code generation scenarios.
III Methodology
In this section, we present the detailed methodology of SEED. Specifically, given a code generation scenario/domain with a limited-sample training dataset $\mathcal{D}_{train}=\{(r,c)\}$, where each sample pair $(r,c)$ consists of an input requirement $r$ and an associated example of desired output code $c$. For a pre-trained LLM $\mathcal{M}_{\theta}$ with parameter $\theta$, we aim to adapt $\mathcal{M}_{\theta}$ to the specific scenario of $\mathcal{D}_{train}$. SEED achieves sample-efficient adaptation of LLMs through four steps: Error Code Collection (§III-A), Automatic Code Revision (§III-B), Model Optimization (§III-C), and Iterative Adaptation (§III-D). The overview of SEED and its differences from traditional fine-tuning are demonstrated in Figure 3.
III-A Error Code Collection
In this step, we systematically identify and collect erroneous output of LLMs using testing as criteria.
We employ rejection sampling [21] to draw error code samples from the distribution produced by $\mathcal{M}_{\theta}$. For each requirement $r\in\mathcal{D}_{train}$, we sample
$c^{\prime}\sim\mathcal{M}_{\theta}(r)~{}|~{}\neg f,$ (1)
where we sample multiple times and employ the criterion function $f$ to determine the retention of the code sample. Specifically, the error code sample $c^{\prime}$ is retained when $f(r,c^{\prime})=0$, where $f(r,c^{\prime})=0$ if the rejection condition is satisfied, otherwise $1$.
We define $f$ as a test evaluation function since testing is the criterion for judging the correctness of code in practice:
$\textsc{TestEval}(r,c^{\prime})=\left\{\begin{array}[]{ll}0,&\text{if }c^{\prime}\text{ fails }S_{r},\\ 1,&\text{otherwise},\end{array}\right.$ (2)
where $S_{r}$ is a suit of test cases under the requirement $r$ and is equipped by the code generation dataset. When collecting the error codes for test failures, we can keep the failed test cases and error messages simultaneously for further error diagnosis.
To gain insights into the propensity of $\mathcal{M}_{\theta}$ to make certain errors, it is advisable to select error code sample $c^{\prime}$ for which the model demonstrates relatively high confidence. Therefore, among multiple error codes collected for the same $r$, we select the one with the highest generation probability  [^3] We determine the probability of the generated code by averaging the probabilities of each generated token..
Figure 4: Illustration of automatic code revision
III-B Automatic Code Revision
In this step, we perform automatic revision for error codes using our Self-Revise approach. Based on the LLM $\mathcal{M}_{\theta}$ itself, Self-Revise revises the error code by providing the information in the original dataset and pointing out the error with code execution feedback. Our objective is to derive a revised code that is as close as possible to the error code. Note that although there is already a correct code $c$ in the dataset, it may differ significantly from the error code, as illustrated by examples (a), (b), and (c) in Figure 3. The pipeline of automatic code revision is shown in Figure 4.
Specifically, we leverage the following parts as the input of Self-Revise:
  • Requirement ($\mathbf{r}$): Clarify the requirement that need to be addressed;
  • Correct Solution ($\mathbf{g}$): Provide a type of correct solution as a reference to reduce the difficulty of revision. The correct solution used here is the code sample $c$ in the training dataset;
  • Error Code ($\mathbf{c^{\prime}}$): Give the error code that needs to be revised. The error code is generated by $\mathcal{M}_{\theta}$ under $r$;
  • Error Messages ($\mathbf{m}$) and Failed Test Cases ($\mathbf{t}$): Point out the error messages received during execution and the specific test cases where the error code fails, allowing for more focused troubleshooting and revision.
These parts are combined as the input of Self-Revise according to the template:
$T=\operatorname{Template}(r,g,c^{\prime},m,t)$ (3)
where $\operatorname{Template}$ is shown in Figure 4.
Following previous work [22, 23], we use two settings for Self-Revise, i.e., fine-tuning (FT) and few-shot prompting (FSP), to get the model $\mathcal{M}_{\operatorname{Revise}}$ for revising error codes.
Self-Revise (FT) entails the process of fine-tuning $\mathcal{M}_{\theta}$ with a small number of samples for the purpose of automatic code revision. The training objective is to minimize the following loss function:
$L(\theta)=\sum_{i}l_{ce}(\mathcal{M}_{\theta}(T_{i}),c^{*}_{i})$ (4)
where $l_{ce}$ represents the standard cross-entropy loss, and we update the parameters initialized with $\mathcal{M}_{\theta}$ to obtain the model $\mathcal{M}_{\operatorname{Revise}}$ in Self-Revise (FT).
Self-Revise (FSP) adopts the few-shot prompting technique and leverages $k$ examples of $T_{i}$ and $c^{*}_{i}$ to construct the prompt $P$ for aligning $\mathcal{M}_{\theta}$ to automatic code revision. In Self-Revise (FSP), $\mathcal{M}_{\operatorname{Revise}}(\mathbf{\cdot})$ is defined as $\mathcal{M}_{\theta}(P\ ||\ \mathbf{\cdot}\ )$, where ||||| | denotes the concatenation operation.
In contrast to the previous error code collection step, for each error code $c^{\prime}$, we construct $T$ and use acceptance sampling to obtain the revised code $c^{*}$:
$c^{*}\sim\mathcal{M}_{\operatorname{Revise}}(T)~{}|~{}f.$ (5)
where $c^{*}$ is retained if $\textsc{TestEval}(r,c^{*})=1$ in Eq. (2), i.e., the revised code $c^{*}$ passes its test cases. We sample multiple times and it is sufficient if $\mathcal{M}_{\operatorname{Revise}}$ could correctly revise the error code once. To prevent $\mathcal{M}_{\operatorname{Revise}}$ from simply replicating the provided correct solution $g$, we exclude the output identical to $g$. Subsequently, for each requirement $r$, and select the version that is most similar to the error code among the remaining code revisions.
III-C Model Optimization
In this step, we employ pairs of the requirement $r$ and its revised code $c^{*}$ to further fine-tune the model $\mathcal{M}_{\theta}$. This process leads to the enhanced version of $\mathcal{M}_{\theta}$, referred to as $\mathcal{M}_{\theta^{*}}$, in the specific scenario of dataset $\mathcal{D}_{train}$.
For fine-tuning [24], we update the entire parameter $\theta$ of LLMs as
$\theta^{*}=\arg\min_{\theta}\sum_{(r,c^{*})}l_{ce}(\mathcal{M}_{\theta}(r),c^{*}),$ (6)
When the computational resources are insufficient, we employ Low-Rank Adaptation (LoRA) [25] to fine-tune LLMs. For a weight matrix $W\in\mathbb{R}^{d\times k}$, LoRA represents its update with a low-rank decomposition:
$W+\Delta W=W+\Delta\alpha W_{\operatorname{down}}W_{\operatorname{up}},$ (7)
where $\alpha$ is a tunable scalar hyperparameter, $W_{\operatorname{down}}\in\mathbb{R}^{d\times r}$, $W_{\operatorname{up}}\in\mathbb{R}^{r\times k}$, and $r<<min(r,k)$. In LoRA, we update parameter $\theta^{*}$ as
$\theta^{*}=\theta+\Delta\theta,$ (8)
$\Delta\theta=\arg\min_{\Delta\theta}\sum_{(r,c^{*})}l_{ce}(\mathcal{M}_{\theta+\Delta\theta}(r),c^{*}).$ (9)
III-D Iterative Adaptation
The preceding three steps can go through multiple iterations until a certain number of rounds is reached or the performance no longer improves.
For $l$-th iteration that $l>1$, we initialize its initial model $\mathcal{M}_{\theta_{l}}$ as the enhanced model of the previous iteration $\mathcal{M}_{\theta^{*}_{l-1}}$. Based on $\mathcal{M}_{\theta_{l}}$, we repeat the process in steps of error code collection and automatic code revision to sample error codes $\{c^{\prime}\}_{l}$ and revised codes $\{c^{*}\}_{l}$, respectively. We use the union of $\{(r,c^{*})\}_{1:l}$, that is,
$\{(r,c^{*})\}_{1}\cup\cdots\cup\{(r,c^{*})\}_{i}\cdots\cup\{(r,c^{*})\}_{l},$ (10)
to update parameters in the model optimization step, thereby yielding the enhanced model of the $l$-th iteration $\mathcal{M}_{\theta^{*}_{l}}$. At each iteration, the model is trained to convergence.
This iteration is a step-by-step optimization designed to continuously improve the adaptability of models to the specific scenario. The complete process of SEED is summarized in Algorithm 1.
Algorithm 1 Pseudocode of the proposed SEED.
0:  Few-sample Dataset $\mathcal{D}_{train}=\{(r,c)\}$, initial LLM $\mathcal{M}_{\theta}$.
0:  LLM $\mathcal{M}_{\theta^{*}}$.
1:  Initial iteration index $l=0$ and $\mathcal{M}_{\theta_{l+1}}=\mathcal{M}_{\theta}$.
2:  # Iterative Adaptation
3:  repeat
4:     Update $l=l+1$.
5:     # Error Code Collection
6:     Perform rejection sampling to collect error codes $\{c^{\prime}\}_{l}$ based on $\mathcal{M}_{\theta_{l}}$ via Eq. (1) and (2).
7:     # Automatic Code Revision
8:     Perform acceptance sampling to collect revised codes $\{c^{*}\}_{l}$ based on $\mathcal{M}_{\theta_{l}}$ and Self-Revise via Eq. (2), (3), and (5).
9:     Calculate the union of $\{(r,c^{*})\}_{1:l}$ via Eq. (10).
10:     # Model Optimization
11:     Fine-tune $\mathcal{M}_{\theta_{l}}$ to yield $\mathcal{M}^{*}_{\theta_{l}}$ via Eq. (6) if the computational resources are sufficient, otherwise via Eq. (7), (8), and (9).
12:     Update $\mathcal{M}_{\theta_{l+1}}=\mathcal{M}_{\theta^{*}_{l}}$.
13:  until End condition is satisfied
14:  return  $\mathcal{M}_{\theta^{*}_{l}}$
IV Evaluation
We present extensive experiments that span three representative code generation datasets, two fine-tuning settings, and five different LLMs of varying sizes or series. We aim to investigate six research questions:
  • RQ1: How does SEED perform compared to the mainstream adaptation approaches?
  • RQ2: How does SEED perform when applied to various LLMs?
  • RQ3: What kind of training sample has the better training effect?
  • RQ4: How does the number of iterations affect the effectiveness of SEED?
  • RQ5: What is the impact of implementing the automatic code revision component of SEED in conjunction with alternative LLMs?
  • RQ6: How does each input component of Self-Revise contribute to the effectiveness?
IV-A Datasets
MBPP [26] contains crowd-sourced Python programming problems, covering programming fundamentals. We selected the version in the work [27], which consists of 276 problems and some generated error codes alongside their human-revised counterparts, thus facilitating subsequent experiments.
HumanEval [1] is a widely-used code generation benchmark, containing 164 handwritten programming problems, proposed by OpenAI. Each programming problem includes a function signature, a natural language description, use cases, a correct solution in Python, and several test tests.
DS-pandas [28] comprises 291 data science problems utilizing Pandas libraries, sourced from real-world problems posted by developers on StackOverflow. This dataset can evaluate the ability of LLMs to utilize specific data-analysis libraries for code generation.
Due to the scarcity and privacy of domain data, it becomes difficult to obtain these data, and there are no available datasets. Therefore, we use the MBPP, HumanEval, and DS-pandas to simulate the specific scenario with limited samples. We choose the first 100 problems from all datasets as test dataset $\mathcal{D}{\text{test}}$. The remaining problems (up to 200 problems) serve as training dataset $\mathcal{D}{\text{train}}$. The statistics of all datasets are presented in Table I.
TABLE I: Statistics of MBPP, HumanEval, and DS-pandas.
Dataset Num of Problems Avg. Length Avg. Num
$\mathcal{D}_{\text{train}}$ $\mathcal{D}_{\text{test}}$ Requirement Code Test Case
MBPP 176 100 15.7 32.5 3
HumanEval 64 100 23.0 30.5 8
DS-pandas 191 100 184.8 26.2 2
IV-B Implementation Details
We use a single A6000 GPU to conduct all experiments. We select CodeGen-2B [5] as our base model by default, which is a well-known open-source LLM for code and is suitable for full fine-tuning within our computational resource constraints. $\mathcal{M}_{\theta}$ is initialized to our base model, and $\mathcal{M}_{\operatorname{Revise}}$ is derived from $\mathcal{M}_{\theta}$ though Self-ReviseIII-B).
For full parameter fine-tuning, i.e., Fine-tuning (Full) [24], we use the AdamW optimizer [29], with hyperparameters $\beta_{1}=0.9$ and $\beta_{2}=0.9$, accompanied by a linear learning rate schedule. The initial learning rate is set to 5e-6, with a batch size of 1 and gradient accumulation of 32 steps for training across 10 epochs. For parameter-efficient fine-tuning, i.e., Fine-tuning (LoRA) [25], the learning rate is set to 2e-4. Additionally, the rank $r$ is adjusted to 128, and the scaling factor $\alpha$ is set at 8. All other hyperparameters remain aligned with Fine-tuning (Full). For few-shot prompting [30], we set the number of examples in prompt to 4. All baselines in the experiments use consistent settings.
In the error code collection step (§III-A) and the automatic code revision step (§III-B), we use temperature [31, 32] sampling to generate multiple samples: 5 samples in the former and 30 in the latter, with the temperature set to 0.8. To obtain the final revised code in the automatic code revision step, we choose the one of revised code exhibiting the minimum Levenshtein distance [33] to the error code. The number of iterative adaptations is set to 2. The maximum generation length is uniformly limited to 256 tokens.
IV-C Metrics
Following the practice of real software development which utilizes testing for evaluation [34, 35], we employ the Pass@k [36] metric to measure the functional correctness of the generated code by executing test cases. We use the unbiased version [1] of Pass@k, where $n>=k$ samples are generated for each problem, count the number of correct samples $c<=n$ which pass test cases and calculate the following estimator,
$\footnotesize{\operatorname{Pass@k}=\mathop{\mathbb{E}}\limits_{\operatorname{Problems}}\begin{bmatrix}1-\frac{\begin{pmatrix}n-c\\ k\end{pmatrix}}{\begin{pmatrix}n\\ k\end{pmatrix}}\end{bmatrix}.}$ (11)
For automatic code revision, we add the pass@any metric which refers to the percentage of tasks for which the model generates at least one correct code that passed all test cases.
In the final evaluation of this paper, we set the temperature to 0.8 and generate $n=50$ samples, which are then used to calculate unbiased Pass@k [1] via Eq. (11) in all experiments. All evaluation results are averaged over five test runs.
IV-D RQ1: Comparison of SEED and mainstream adaptation approaches
Baselines. In this section, we evaluate the effectiveness of SEED by comparing it against other four mainstream approaches used as baselines:
  • Direct Generation: We evaluate the LLM directly to demonstrate the performance of the original LLM.
  • Fine-tuning (Full): We employ full-parameter fine-tuning for the LLM on $\mathcal{D}_{train}$.
  • Fine-tuning (LoRA): We fine-tune the LLM on $\mathcal{D}_{train}$ using LoRA [25], which greatly reduces the number of trainable parameters.
  • Few-shot Prompting: We use 4-shot prompt [30] to align LLMs with the input-output format of $\mathcal{D}_{train}$, where 4 examples in prompt are randomly selected from $\mathcal{D}_{train}$.
Considering the baselines involve full-parameter fine-tuning, CodeGen-2B is uniformly selected as the base model in this experiment. For SEED, we use 30% data of $\mathcal{D}_{train}$ for Self-Revise (FT) [^4] In addition to MBPP dataset, for other two datasets (i.e., HumanEval and DS-pandas), we generate one error code per sample in a subset comprising 30% of the training set, using CodeGen-2B. Subsequently, authors collaboratively apply the minimal necessary revisions to correct these error codes., while the remaining 70% data of $\mathcal{D}_{train}$ is employed for model optimizing, where we use full-parameter fine-tuning.
TABLE II: Pass@k (%) of SEED and baselines on MBPP, HumanEval, and DS-pandas benchmarks, and the teal number after $\uparrow$ denotes the relative improvement of SEED over Fine-tuning (Full).
Benchmark Approach Pass@1 Pass@5 Pass@10
MBPP Direct Generation 15.6% 31.4% 40.2%
Fine-tuning (Full) 25.8% 46.2% 57.6%
Fine-tuning (LoRA) 19.8% 39.8% 55.2%
Few-shot Prompting 24.4% 38.0% 49.4%
[…]
HumanEval Direct Generation 24.8% 44.7% 51.8%
Fine-tuning (Full) 29.8% 47.9% 56.4%
Fine-tuning (LoRA) 27.4% 46.9% 53.9%
Few-shot Prompting 25.2% 45.8% 53.1%
[…]
DS-pandas Direct Generation 0.8% 3.1% 5.6%
Fine-tuning (Full) 2.6% 6.5% 9.6%
Fine-tuning (LoRA) 2.2% 6.0% 8.9%
Few-shot Prompting 1.9% 4.5% 5.7%
[…]
Results. We conducted experiments on three public benchmarks, i.e., MBPP, HumanEval, and DS-pandas. The experimental results are summarized in Table IV-D. This comparison yielded three insightful observations: 1) Significant superiority of SEED: Our proposed SEED performs significantly better than the other four baselines on the three benchmarks. Notably, SEED exhibits significant relative improvements of 27.2%, 32.9%, and 103.9%, respectively, when compared to the best-performing baseline Fine-tuning (Full). 2) Worst performance of Direct Generation: The performance of Direct Generation is significantly lower than the Fine-tuning (Full), Fine-tuning (LoRA), and Prompt baselines. This result suggests that directly applying LLMs for evaluation may be less suitable for specific scenarios, resulting in performance differences. 3) Fine-tuning (LoRA) is less effective than Fine-tuning (Full): Although Fine-tuning (LoRA) offers the advantage of reduced computational resource requirements for fine-tuning LLMs, it trades off the performance. 4) Less improvement of Few-shot Prompting: Few-shot Prompting generally underperforms compared to the Fine-tuning approaches. It only outperforms Fine-tuning (LoRA) in the MBPP benchmark’s Pass@1 metric. However, in other situations, the efficacy of Few-shot Prompting is the lowest, with the exception of Direct Generation.
Summary of RQ1:  In code generation scenarios with limited training samples, SEED exhibits improvements compared to the mainstream adaptation approaches, achieving relative improvements between 27.2% and 103.9%.